A 1 ohm resistor is connected in parallel to a 10 ohm resistor. Calculate the equivalent resistance.

Grade:12

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
In a parallel combination of resistors,
The equivalent resistance is given as
\frac{1}{R_{eq}}=\frac{1}{R_{1}}+\frac{1}{R_{2}}
So R1 is 1 and R2 is 10 ohms. and hence the net resistance becomes 10/11 ohms.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty